# Hard View Selection for Self-Supervised Learning

### v4

@@ Abstract
Many Self-Supervised Learning (SSL) methods train their models to be invariant to different “views” of an image and considerable efforts were directed towards improving pre-text tasks, architectures, or robustness. However, most SSL methods remain reliant on the random sampling of operations within the image augmentation pipeline, such as the random resized crop operation. We argue that the role of the view generation and its effect on performance has so far received insufficient attention. To address this, we propose an easy, learning-free, yet powerful Hard View Selection (HVS) strategy designed to extend the random view generation to expose the pretrained model to harder samples during SSL training. It encompasses the following iterative steps: 1) randomly sample multiple views and create pairs of two views, 2) run forward passes for each view pair on the currently trained model, 3) adversarially select the pair yielding the worst loss depending on the current model state, and 4) run the backward pass with the selected pair. As a result, HVS consistently achieves accuracy improvements between 0.91% and 1.93% on ImageNet linear evaluation and similar improvements on transfer tasks across DINO, SimSiam, iBOT and SimCLR. We provide studies to shed light on the inner workings and show that, by naively using smaller resolution images for the selection step, we can significantly reduce the computational overhead while retaining performance. Surprisingly, even when accounting for the computational overhead incurred by HVS, we achieve performance gains between 0.52% and 1.02% and closely rival the 800-epoch DINO pretraining with only 300 epochs.
@@ 1 Introduction
Figure 1: HVS first samples $N$ views, pairs them, and adversarially selects the hardest pair, i.e., the one with the worst loss according to the current model state.
@@
Various approaches to learn effective and generalizable visual representations in Self-Supervised Learning (SSL) exist. One way to categorize many SSL methods is to distinguish generative and discriminative approaches. While generative methods aim at generating image input, discriminative methods, with contrastive learning (Hadsell et al., 2006; He et al., 2020) being the most renowned approach, aim at learning a latent representation in which similar image views are closely and dissimilar ones are distantly located.
@@
Such views are generated by applying a sequence of (randomly sampled) image transformations and are usually composed of geometric (cropping, rotation, etc.) and appearance (color distortion, blurring, etc.) transformations. A body of literature (Chen et al., 2020a; Wu et al., 2020; Purushwalkam & Gupta, 2020; Wagner et al., 2022; Tian et al., 2020b) has illuminated the effects of image views on representation learning and identified random resized crop (RRC) transformation, which randomly crops the image and resizes it back to a fixed size, as well as color distortion as one of the most essential ones for effective contrastive learning. However, despite this finding and to our best knowledge, little research has gone into identifying more effective ways for selecting or generating views to improve performance. In this paper, we seek to extend the random view generation to expose the pretrained model to harder samples during CL training.
@@
Figure 2: Examples (left) and sampled views (right) after resizing and appearance transformations. Pairs selected by HVS are shown with a solid frame.
Supervised Learning works that address this shortcoming and which attempt to leverage task difficulty include Spatial Transformer Networks (Jaderberg et al., 2015) and Adversarial AutoAugment (Zhang et al., 2020), which either directly learn geometric transformations or generate augmentation policies that yield hard samples. However, these approaches have not been used successfully for unsupervised learning and Adversarial AutoAugment does not include RRC in its search space. In contrast to what we propose, integrating any of these methods into existing self-supervised learning pipelines is non-trivial.
Related SSL approaches exist that aim at controlling view hardness through adversarial Shi et al. (2022); Tamkin et al. (2021) or cooperative Hou et al. (2023) models. A work that takes a similar perspective like us in leveraging the information content of views for SSL is (Tian et al., 2020b). They discover sweet spots for data augmentation hyperparameters that are widely used and adversarially learn flow-based models that generate novel color space-views which are then used for contrastive learning. While these works offer important insights into view optimality, they usually require learning additional models prior to the actual SSL training or non-trivial changes to the pipeline and thus, are not easily integrated into existing SSL methods.
In contrast, we propose a fully learning-free and easy-to-integrate approach. Importantly, and different to Tian et al. (2020b), our approach leverages the current model state and operates on a per-sample basis during pretraining. It applies the following simple strategy: we sample a few views and forward each pair through the currently trained model. Next, we identify the pair that yields the worst loss among all pairs, use it for the backward pass, and repeat (see Fig. 1). While being agnostic to any contrastive and non-contrastive loss, this strategy exposes the pretrained model to samples that the model at any given point in the training trajectory finds challenging. As our results show, sampling only four views are sufficient to achieve noticeable improvements.
@@
  • We propose Hard View Selection (HVS), a new learning method complementing SSL that is easy to integrate and extends the broadly-used random view generation to automatically expose the model to harder samples during pretraining. The only requirement of HVS is that it needs to be able to compute sample-wise losses;
  • We demonstrate the effectiveness and compatibility of our approach using ImageNet (Deng et al., 2009) pretraining across popular and widely-adopted SSL methods that cover a diverse range of objectives such as SimSiam (Chen & He, 2021), DINO (Caron et al., 2021), iBOT (Zhou et al., 2021), and SimCLR (Chen et al., 2020a);
  • HVS improves all baselines noticeably compared to using the common view generation, and we show this on a diverse set of transfer datasets and tasks, including linear evaluation, finetuning, object detection and segmentation;
  • We present insights into the underlying mechanisms of HVS such as the effect of changing the number of sampled views or color augmentation.
We make our PyTorch (Paszke et al., 2019) code, models, and all used hyperparameters publicly available under https://anonymous.4open.science/r/hard-view-selection/.
@@ 2 Related Work
2.1 Contrastive and Non-Contrastive Learning in SSL
The core idea behind the contrastive learning framework is to learn image representations by contrasting positive pairs (two views of the same image) against negative pairs (two views of different images) (Hadsell et al., 2006). To work well in practice and to prevent model collapse, contrastive learning methods often require a large number of negative samples (Wu et al., 2018; van den Oord et al., 2018; Chen et al., 2020a; He et al., 2020; Tian et al., 2020a; Chen et al., 2020b) stored in memory banks (Wu et al., 2018; He et al., 2020) or, for instance, in the case of SimCLR, implicitly in large batches (Chen et al., 2020a). Non-contrastive approaches, such as BYOL (Grill et al., 2020), SimSiam (Chen & He, 2021), DINO (Caron et al., 2021) and others (Zbontar et al., 2021; Caron et al., 2020; Ermolov et al., 2021), are able to only use positive pairs without causing model collapse but rely on other techniques, such as Siamese architectures, whitening of embeddings, clustering, maximizing the entropy of the embeddings, momentum encoders, and more.
2.2 Optimizing for Hard Augmentations and Views in SSL
Past work has shown that optimizing augmentation policies or views directly can be beneficial. Hence, the realm of learning task-specific augmentation policies based on data has seen quick development (Cubuk et al., 2019; Ho et al., 2019; Lin et al., 2019; Zhang et al., 2020; Hataya et al., 2020; Müller & Hutter, 2021). Most related to our work among these approaches is Adversarial AutoAugment (AAA) (Zhang et al., 2020) that proposes to generate augmentation policies yielding hard samples by learning with an adversarial supervised objective. In a similar, but cooperative setting, MADAug (Hou et al., 2023) learns policies yielding progressively harder samples. In contrast to our work, however, AAA and MADAug do not include the random resize crop operation into their search spaces and thus, are limited in leveraging the geometric perspective to control view hardness.
Unlike learning hard augmentations, approaches exist that optimize directly on the pixel level to control view hardness. For instance, Shi et al. (2022) adversarially learn semantically meaningful masks for a pretrained network to inpaint. Another approach perturbs input images using an adversarial model (Tamkin et al., 2021). Tian et al. (2020b) leverage, like us, the view content for better performance. Under the lens of mutual information (MI), they derive an objective to adversarially learn flow-based view generators for contrastive learning. While these methods can complement our approach, they are less easily integrated into existing contrastive pipelines due to the need for adversarial network training. Moreover, Tian et al. (2020b) does not use the current model state to control learning hardness and, while their studies incorporate distance metrics, they do not consider the Intersection over Union (IoU) metric which we extensively make use of.
@@ 3 Method
3.1 Contrastive Learning Framework
In this section, we introduce our approach which is also depicted in Algorithm 12. Many different contrastive learning (He et al., 2020) objectives exist, each characterized by variations stemming from design choices, such as by the use of positive and negative samples or asymmetry in the encoder-projector network structure. For simplicity of exposure, we will introduce our approach based on the SimSiam objective Chen & He (2021), but we do note that our method can be used with any other discriminative SSL objective that allows the computation of a sample-wise losses.
@@ 3.1 Contrastive Learning Framework
3.2 Hard View Selection
As previously described, Hard View Selection extends the random view generation by sampling adversarially harder views during pretraining. Instead of having two sets of augmentations $A$ and $B$, we now sample $N$ sets of augmentations, denoted as $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{N}\}$. Each set $A_{i}$ is sampled from the image augmentation distribution $\mathcal{T}$, and applied to each image in $\mathbf{x}$, resulting in $N$ augmented sets of views $\mathbf{x}^{A_{1}},\mathbf{x}^{A_{2}},\ldots,\mathbf{x}^{A_{N}}$. Similarly, we obtain $N$ sets of embeddings $\mathbf{z}^{A_{1}},\mathbf{z}^{A_{2}},\ldots,\mathbf{z}^{A_{N}}$ and predictions $\mathbf{h}^{A_{1}},\mathbf{h}^{A_{2}},\ldots,\mathbf{h}^{A_{N}}$ through the encoder and projector networks. We then define a new objective function that seeks to find the pair $(x_{i}^{A_{k}},x_{i}^{A_{l}})$ of a given image $x_{i}$ that yields the highest loss:
$(x_{i}^{A_{k}},x_{i}^{A_{l}})=\operatorname*{arg\,max}_{k,l;k\neq l}\mathcal{L}(\theta)_{i,k,l}=\operatorname*{arg\,max}_{k,l;k\neq l}\frac{1}{2}\left(D(z_{i}^{A_{k}},h_{i}^{A_{l}})+D(z_{i}^{A_{l}},h_{i}^{A_{k}})\right),$ (2)
where $\mathcal{L}(\theta)_{i,k,l}$ simply denotes a sample-wise variant of Eq. 1. Overall, we first generate $N$ augmented views for each image $x_{i}$ in the minibatch. Then, we create all combinatorially possible $\binom{N}{2}$ pairs of augmented images for $x_{i}$. Then, we use Eq. 2 to compute the sample-wise loss for each pair which corresponds to invoking forward passes. We then select all pairs that yielded the highest loss to form the new hard minibatch of augmented sets $\textbf{x}^{A_{k_{*}}}$ and $\textbf{x}^{B_{l_{*}}}$, discard the other pairs and use the hard minibatch for optimization.
@@ 3.2 Hard View Selection
2:augmentation distribution $\mathcal{T}$, encoder $f_{\theta}$, projector $g_{\theta}$
3:for each $x_{i}$ in the sampled minibatch $\left\{x_{i}\right\}_{i=1}^{M}$ do
@@
6:     Forward all views through $f_{\theta}$ and $g_{\theta}$ to get $z$ and $h$
7:     Create all $\binom{N}{2}$ possible pairs $\textbf{x}_{i}^{A_{k}}\times\textbf{x}_{i}^{A_{l}}$, $k\neq l$
8:     Select the hard pair $(x_{i}^{A_{k_{*}}},x_{i}^{A_{l_{*}}})$ maximizing Eq. 2
9:Create the minibatch $(\textbf{x}^{A_{k_{*}}},\textbf{x}^{A_{l_{*}}})$ consisting of hard pairs
10:Proceed with standard contrastive learning training
to update $\theta$ with the hard batch
11:Repeat for all minibatches
12:return Optimized $\theta$
Algorithm 1 Contrastive Learning with HVS
As shown in Algorithm 12, we repeat this process in each training iteration. This modification introduces a more challenging learning scenario in which the model is encouraged to learn more discriminative features by being exposed to harder views. While the additional forward passes come with a certain overhead (about a factor of 1.55$\times$ for SimSiam), as we will demonstrate in our experiments, the focus on harder views leads to consistent improvements.
@@
While we exemplified the integration of HVS with the SimSiam objective, integrating it into other contrastive methods is as straightforward. The only requirement of HVS is to be able to compute sample-wise losses (in order to select the views with the highest loss). In our experiments section and in addition to SimSiam, we study the application of HVS to the objectives of DINO and SimCLR (see also Appendix I.1 for a formal exemplary definition for the integration of HVS into SimCLR).
3.3 Implementation and Evaluation Protocols
Implementation
@@ 3.3 Implementation and Evaluation Protocols
We now describe the technical details of our approach. HVS can be used with any contrastive method that allows computing sample-wise losses, and the only two elements in the pipeline we adapt are: 1) the data loader (needs to sample $N$ views for each image) and 2) the forward pass (invokes the select function to identify and return the hard views). The image transformation distribution $\mathcal{T}$ of the baselines is left unchanged. Note, for the view selection one could simply resort to RRC only and apply the rest of operations after the hard view selection (see Section 5.1 for a study on the influence of appearance on the selection). All experiments were conducted with $N=4$ sampled views, yielding $\binom{N}{2}=6$ pairs to compare, except for DINO that uses 10 views (2 global, 8 local heads) per default. For DINO, we apply HVS to both global and local heads but to remain tractable, we upper-bound the number of total pair comparisons to 128. SimCLR uses positive and negative samples. In accordance with the simplicity of HVS, we do not alter its objective, which naturally leads to selecting hard views that are adversarial to positive and “cooperative” to negative views.
@@
+ HVS RN50 68.98 58.97 70.90 62.97
Improvement +0.78 +1.50 +0.55 +1.57
SimCLR RN50 63.40 52.83 65.50 55.65
+ HVS RN50 65.33 54.76 67.20 57.31
Improvement +1.93 +1.85 +1.70 +1.66
iBOT ViT-S/16 69.50 62.82 72.48 66.60
+ HVS ViT-S/16 70.41 63.05 73.99 67.12
Improvement +0.91 +0.23 +1.51 +0.52
Table 1: Average top-1 linear and k-NN classification accuracy on the ImageNet validation set for 100 and 300-epoch pretrainings. For SimCLR we run 200 instead of 300 epochs.
Evaluation Protocols

### v6

# Beyond Random Augmentations: Pretraining with Hard Views

@@ Abstract
Self-Supervised Learning (SSL) methods typically rely on random image augmentations, or views, to make models invariant to different transformations. We hypothesize that the efficacy of pretraining pipelines based on conventional random view sampling can be enhanced by explicitly selecting views that benefit the learning progress. A simple yet effective approach is to select hard views that yield a higher loss. In this paper, we propose Hard View Pretraining (HVP), a learning-free strategy that extends random view generation by exposing models to more challenging samples during SSL pretraining. HVP encompasses the following iterative steps: 1) randomly sample multiple views and forward each view through the pretrained model, 2) create pairs of two views and compute their loss, 3) adversarially select the pair yielding the highest loss according to the current model state, and 4) perform a backward pass with the selected pair. In contrast to existing hard view literature, we are the first to demonstrate hard view pretraining’s effectiveness at scale, particularly training on the full ImageNet-1k dataset, and evaluating across multiple SSL methods, ConvNets, and ViTs. As a result, HVP sets a new state-of-the-art on DINO ViT-B/16, reaching 78.8% linear evaluation accuracy (a 0.6% improvement) and consistent gains of 1% for both 100 and 300 epoch pretraining, with similar improvements across transfer tasks in DINO, SimSiam, iBOT, and SimCLR.
@@ 1 Introduction
Learning effective and generalizable visual representations in Self-Supervised Learning (SSL) has been approached in various ways. Many SSL methods can be broadly categorized into generative and discriminative approaches (Chen et al., 2020a). Generative methods focus on generating image inputs, while discriminative methods, particularly contrastive learning (Hadsell et al., 2006; He et al., 2020), aim at learning latent representations in which similar image views are located closely, and dissimilar ones distantly.
@@
Such views are generated by applying a sequence of (randomly sampled) image transformations and are usually composed of geometric (cropping, rotation, etc.) and appearance (color distortion, blurring, etc.) transformations. Prior work (Chen et al., 2020a; Wu et al., 2020; Purushwalkam & Gupta, 2020; Wagner et al., 2022; Tian et al., 2020b) has identified random resized crop (RRC), which randomly crops the image and resizes it back to a fixed size, as well as color distortion as critical transformations for effective representation learning. However, despite this finding and to our best knowledge, little research has gone into identifying more effective ways for generating views to improve performance.
@@
Figure 1: (a) HVP first samples $N$ views, pairs them, and adversarially selects the hardest pair, i.e., the one with the worst loss according to the current model state. (b) Examples (left) and sampled views (right) after transformations. Hard pairs selected by HVP are shown with a solid frame.
Existing SSL approaches that attempt to control the hardness of views include adversarial (Shi et al., 2022; Tamkin et al., 2021) or cooperative (Hou et al., 2023) techniques. For instance, Tian et al. (2020b) use mutual information theory to adversarially learn view generators. These methods offer valuable insights into optimizing views but often introduce additional complexity, such as requiring additional models or significant changes to the SSL pipeline, limiting their practicality in state-of-the-art models where resource constraints are already a concern. Similar to what we propose, Koçyigit et al. (2023) introduce a hard view sampling strategy but targeting the acceleration of pretraining. Moreover, their approach requires tuning hyperparameters like learning rate and augmentation magnitude, which can introduce confounding variables. Furthermore, neither Koçyigit et al. (2023) nor Tian et al. (2020b) validate their methods on larger datasets such as ImageNet-1k (Deng et al., 2009), or across larger model architectures, limiting their scalability and applicability.
Building on these observations, we propose Hard View Pretraining (HVP), a fully learning-free, easy-to-integrate approach designed to improve standard pretraining methods without the need for additional model training or complex modifications. Our method leverages the current model state to select challenging samples during pretraining by adversarially sampling pairs of views and selecting the pair that yields the highest loss according to the model‘s current state (see Fig. 1(a)). Unlike previous approaches, HVP requires no hyperparameter tuning and demonstrates scalability to large datasets like ImageNet-1k, offering an efficient and practical solution for improving SSL pretraining. To the best of our knowledge, we are the first to demonstrate the effectiveness of a hard view sampling strategy at scale, particularly on modern architectures like Vision Transformers (ViTs) and training on the full ImageNet dataset. Our approach not only integrates seamlessly with recent state-of-the-art SSL methods but also showcases consistent improvements across both convolutional architectures and ViTs, validating its robustness and scalability.
@@
  • We propose Hard View Pretraining (HVP), an easy-to-use method complementing SSL by extending the common random view generation to automatically expose the model to harder samples during pretraining. HVP simply requires the ability to compute sample-wise losses;
  • We demonstrate the effectiveness and compatibility of our approach using ImageNet-1k pretraining across four popular SSL methods that cover a diverse range of discriminative objectives such as SimSiam (Chen & He, 2021), DINO (Caron et al., 2021), iBOT (Zhou et al., 2021), and SimCLR (Chen et al., 2020a);
  • HVP achieves a new state-of-the-art result on DINO ViT-B/16, improving over the officially reported baseline of 78.2% linear evaluation accuracy by reaching 78.8% (400 epochs). HVP also consistently improves all other baselines by an average of 1% in linear evaluation on ImageNet across 100 and 300 epoch-pretraining runs;
  • We show similar improvements on a diverse set of transfer tasks, including finetuning, object detection, and segmentation, and present insights into the underlying mechanisms and robustness of HVP.
We make our PyTorch Paszke et al. (2019) code, models, and all used hyperparameters publicly available under https://github.com/automl/hvp.
@@ 2 Related Work
2.1 Discriminative Self-Supervised Learning
The core idea behind the discriminative learning framework (Chen et al., 2020a) is to learn image representations by contrasting positive pairs (two views of the same image) against negative pairs (two views of different images) (Hadsell et al., 2006). To work well in practice and to prevent model collapse, contrastive learning methods often require a large number of negative samples (Wu et al., 2018; van den Oord et al., 2018; Chen et al., 2020a; He et al., 2020; Tian et al., 2020a; Chen et al., 2020b) stored in memory banks (Wu et al., 2018; He et al., 2020) or, for instance, in the case of SimCLR, implicitly in large batches (Chen et al., 2020a). Non-contrastive approaches, such as BYOL (Grill et al., 2020), SimSiam (Chen & He, 2021), DINO (Caron et al., 2021) and others (Zbontar et al., 2021; Caron et al., 2020; Ermolov et al., 2021), can only use positive pairs without causing model collapse but rely on other techniques, such as Siamese architectures, whitening of embeddings, clustering, maximizing the entropy of the embeddings, momentum encoders, and more.
2.2 Optimizing for Hard Views in SSL
Due to its performance-improving benefits, the realm of learning task-specific augmentation policies based on data has seen quick development (Cubuk et al., 2019; Ho et al., 2019; Lin et al., 2019; Zhang et al., 2020; Hataya et al., 2020; Hou et al., 2023; Müller & Hutter, 2021). However, these approaches do not include the random resize crop operation in their search spaces, limiting the control of view hardness. Similar to us, Koçyigit et al. (2023) uses the current model state for selecting hard views. However, their approach requires controlling learning hyperparameters, while mostly training on a smaller version of ImageNet and ResNets (He et al., 2016) only. We offer a more complete analysis of hard view pretraining, demonstrating the benefits without potential confounding factors such as hyperparameter adjustments. We also focus on performance rather than pretraining speedups and employ higher budgets (longer pretraining, larger batch sizes) on the full ImageNet dataset and both ResNets and ViTs (Dosovitskiy et al., 2020)). Other works utilize additional networks for view generation, such as Tamkin et al. (2021); Shi et al. (2022); Tian et al. (2020b) (adversarial view generators), Peng et al. (2022) (localization network for semantic awareness), Li et al. (2024) (pretrained generative models to enhance augmentation quality), and Han et al. (2023) (leveraging a pretrained GAN in a SimCLR-only setting). However, unlike HVP, all these methods add non-trivial complexity to the training pipeline by requiring learning auxiliary and adversarial components or are often limited in their applicability across different SSL frameworks (e.g., by requiring negative view pairs).
@@ 3 Method
3.1 Self-supervised Learning Framework
In this section, we introduce our approach, which is also depicted in Algorithm 12. Many different self-supervised discriminative learning (He et al., 2020) objectives exist, each characterized by variations stemming from design choices, such as by the use of positive and negative samples or asymmetry in the encoder-projector network structure. For simplicity of exposure, we will introduce our approach based on the SimSiam objective (Chen & He, 2021), but we do note that our method can be used with any other discriminative SSL objective that allows the computation of sample-wise losses.
@@ 3.1 Self-supervised Learning Framework
3.2 Pretraining with Hard Views
We now formalize how we expose the model to more challenging views during pretraining. In a nutshell, Hard View Pretraining extends the random view generation by sampling adversarially harder views during pretraining. Instead of having two sets of augmentations $A$ and $B$, we now sample $N$ sets of augmentations, denoted as $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{N}\}$. Each set $A_{i}$ is sampled from the image augmentation distribution $\mathcal{T}$, and applied to each image in $\mathbf{x}$, resulting in $N$ augmented sets of views $\mathbf{x}^{A_{1}},\mathbf{x}^{A_{2}},\ldots,\mathbf{x}^{A_{N}}$. Similarly, we obtain $N$ sets of embeddings $\mathbf{z}^{A_{1}},\mathbf{z}^{A_{2}},\ldots,\mathbf{z}^{A_{N}}$ and predictions $\mathbf{h}^{A_{1}},\mathbf{h}^{A_{2}},\ldots,\mathbf{h}^{A_{N}}$ through the encoder and projector networks. We then define a new objective function that seeks to find the pair $(x_{i}^{A_{k}},x_{i}^{A_{l}})$ of a given image $x_{i}$ that yields the highest loss:
$(x_{i}^{A_{k}},x_{i}^{A_{l}})$ $=\operatorname*{arg\,max}_{k,l;k\neq l}\mathcal{L}(\theta)_{i,k,l}$ (2)
$=\operatorname*{arg\,max}_{k,l;k\neq l}\frac{1}{2}\left(D(z_{i}^{A_{k}},h_{i}^{A_{l}})+D(z_{i}^{A_{l}},h_{i}^{A_{k}})\right),$
where $\mathcal{L}(\theta)_{i,k,l}$ simply denotes a sample-wise variant of Eq. 1.
@@ 3.2 Pretraining with Hard Views
2:augmentation distribution $\mathcal{T}$, model $f$
3:for each $x_{i}$ in the sampled batch $\left\{x_{i}\right\}_{i=0}^{M}$ do
@@
6:     Forward all views through $f$
7:     Create all $\binom{N}{2}$ view pairs $\textbf{x}_{i}^{A_{k}}\times\textbf{x}_{i}^{A_{l}}$, $k\neq l$
8:     Add hard pair $(x_{i}^{A_{k_{*}}},x_{i}^{A_{l_{*}}})$ that maximizes
9:    Eq. 2 to the new batch with only hard pairs
10:Proceed with standard SSL training
11:Repeat for all batches
12:return Pretrained model $f$
Algorithm 1 Pretraining with Hard Views
Overall, we first generate $N$ augmented views for each image $x_{i}$ in the minibatch. Then, we forward these augmented views through the networks and create all combinatorially possible $\binom{N}{2}$ pairs of augmented images. Subsequently, we use Eq. 2 to compute the sample-wise loss for each pair. We then select all pairs that yielded the highest loss to form the new hard minibatch of augmented sets $\textbf{x}^{A_{k_{*}}}$ and $\textbf{x}^{B_{l_{*}}}$, discard the other pairs and use the hard minibatch for optimization. As shown in Algorithm 12, we repeat this process in each training iteration.
Intuitively, our approach introduces a more challenging learning scenario in which the model is encouraged to learn more discriminate features by being exposed to harder views. In the early stage of training, the embedding space lacks a defined structure for representing similarity among views. As training progresses, our method refines the concept of similarity through exposure to views that, from the perspective of the model, remain challenging given its current state. By limiting the number of sampled views, we upper-bound the difficulty of learning to prevent tasks from becoming too difficult and hindering learning. This ensures a controlled evolution of the embedding space, where the model’s perception of difficulty is continuously challenged in tandem with its growing capacity to differentiate views. Consequently, HVP can be seen as a regularization that prevents the model from overfitting to easy views.
@@
While we exemplified the integration of HVP with the SimSiam objective, integrating it into other contrastive methods is as straightforward. The only requirement of HVP is to be able to compute sample-wise losses (to select the views with the highest loss). In our experiments section and in addition to SimSiam, we study the application of HVP to the objectives of DINO, iBOT, and SimCLR (see also Appendix K.1 for a formal exemplary definition for the integration of HVP into SimCLR).
4 Implementation and Evaluation Protocols
4.1 Implementation
@@
We now describe the technical details of our approach. HVP can be used with any SSL method that allows computing sample-wise losses, and the only two elements in the pipeline we adapt are: 1) the data loader (which now needs to sample $N$ views for each image) and 2) the forward pass (which now invokes a select function to identify and return the hard views). The image transformation distribution $\mathcal{T}$ taken from the baselines is left unchanged. Note, for the view selection one could simply resort to random resized crop (RRC) only and apply the rest of the operations after the hard view selection (see Section 6.1 for a study on the influence of appearance on the selection).
All experiments were conducted with $N=4$ sampled views, yielding $\binom{N}{2}=6$ pairs to compare, except for DINO which uses 10 views (2 global, 8 local heads) per default. For DINO, we apply HVP to both global and local heads but to remain tractable, we upper-bound the number of total pair comparisons to 128. SimCLR uses positive and negative samples. Following the simplicity of HVP, we do not alter its objective, which naturally leads to selecting hard views that are adversarial to positive and “cooperative" to negative views. For iBOT, we use the original objective as defined by the authors with global views only.
@@
+ HVP RN50 68.98 58.97 70.90 62.97
+0.78 +1.50 +0.55 +1.57
SimCLR RN50 63.37 52.83 65.50 55.65
+ HVP RN50 65.33 54.76 67.30 56.80
+1.96 +1.93 +1.80 +1.15
iBOT ViT-S/16 69.55 62.93 72.76 66.92
+ HVP ViT-S/16 70.27 62.75 73.99 67.16
+0.73 -0.18 +1.23 +0.24
Table 1: Average top-1 linear and k-NN classification accuracy on the ImageNet validation set for 100 and 300-epoch pretrainings across 3 seeds.
4.2 Evaluation Protocols